Cued by celebrities such as Serena Williams and Michelle Obama’s candid discussions of pregnancy-related complications, in recent months, maternal mortality and adverse birth outcomes in the United States have become a hot topic of interest. New review committees have been formed and interventions developed because the United States has the highest rates of maternal mortality of any developed nation. One such method of reducing the incidence of maternal mortality is to ensure women have continuous prenatal care. Say for example, a new intervention that aims to increase the number of San Francisco women getting regular prenatal screenings is established. Before this program was instituted, a certain area of San Francisco reported that 21% of pregnant women did not have regular prenatal appointments. After the interventions was implemented, a follow-up was conducted with 175 pregnant women. Of those women, 15% reported that they did not have regular prenatal appointments. Does this result suggest that the intervention is effective?

  1. No, because the standard deviation = 3.78, which is less than the population mean, the intervention was not effective.
  2. Yes, because the sample mean = 26.25, which is less than the population mean 36.75, the program is effective.
  3. No, because the number of women who reported no regular prenatal appointments falls in the interval based on the range rule of thumb, the program is not effective.
  4. Yes, because the number of women who reported no regular prenatal appointments falls outside the interval based on the range rule of thumb, the program is effective.
  5. The effectiveness of the program cannot be determined with the given information.
Answer #1
here population proportion=     p= 0.210
sample size       =n= 175
std error of proportion=σp=√(p*(1-p)/n)= 0.0308
minimum usual value =μ-2*σ = 0.1484
maximum usual value =μ-2*σ = 0.2716

option C is correct No, because the number of women who reported no regular prenatal appointments falls in the interval based on the range rule of thumb, the program is not effective.